# Leveraging Large Language Models to Measure Gender Bias in Gendered Languages

### v1

@@ Abstract
Gender bias in text corpora used in various natural language processing (NLP) contexts, such as for training large language models (LLMs), can lead to the perpetuation and amplification of societal inequalities. This is particularly pronounced in gendered languages like Spanish or French, where grammatical structures inherently encode gender, making the bias analysis more challenging. Existing methods designed for English are inadequate for this task due to the intrinsic linguistic differences between English and gendered languages. This paper introduces a novel methodology that leverages the contextual understanding capabilities of LLMs to quantitatively analyze gender representation in Spanish corpora. By utilizing LLMs to identify and classify gendered nouns and pronouns in relation to their reference to human entities, our approach provides a nuanced analysis of gender biases. We empirically validate our method on four widely-used benchmark datasets, uncovering significant gender disparities with a male-to-female ratio ranging from 4:1 to 6:1. These findings demonstrate the value of our methodology for bias quantification in gendered languages and suggest its application in NLP, contributing to the development of more equitable language technologies.
Leveraging Large Language Models to Measure
Gender Bias in Gendered Languages
@@ 1 Introduction
Figure 1: Overview of the proposed method for gender bias detection and measurement in gendered languages using LLMs.
@@
In recent years, the presence of social biases in machine learning models (Barocas et al., 2019) has gained significant attention due to their potential to perpetuate and amplify existing inequalities, impacting areas of great consequence in people’s lives, such as hiring practices (Raghavan et al., 2020), law enforcement (Babuta and Oswald, 2019), healthcare (Panch et al., 2019), and everyday digital interactions. Among various forms of bias, gender bias is particularly concerning because it affects roughly half of the global population and has pervasive effects across different sectors of society. This concern is amplified in the area of natural language processing (NLP), where biases present in training corpora can propagate through machine learning models and impact a broad range of applications.
@@
In this work, we consider bias in text corpora used to train large language models (LLMs). The main motivation for studying bias in training data of LLMs stems from the lifecycle of these models. From training to deployment, biases in the training data can not only be perpetuated but even amplified by the models (Kotek et al., 2023). This data often comes from diverse sources, including books, websites, social media, and other digital texts. If these sources contain biases – which they frequently do, reflecting societal prejudices and stereotypes – the models learn these biases as part of their training process. Hence, the first stage – data curation – becomes of paramount importance. It involves selecting, cleaning, and annotating the training data to minimize biases as much as possible. Effective data curation requires a nuanced understanding of the types of biases that can exist and the methods to mitigate them, including removing or balancing biased content, ensuring diverse and representative data samples, and incorporating fairness and ethical guidelines in the data selection process (Gallegos et al., 2024).
However, existing methods for evaluating bias in text, such as gender polarity (Dhamala et al., 2021) and co-occurrences (Bordia and Bowman, 2019), have been primarily developed for English. These approaches often rely on the neutrality of the English language in terms of grammatical gender, allowing for relatively straightforward identification and categorization of gendered expressions. Nonetheless, such an approach is not directly applicable to gendered languages like Spanish, where nouns, articles, adjectives and other word forms reflect gender, making the detection and analysis of gender bias a much harder problem to solve.
Given the limitations of existing methods and the prevalence of gendered languages in the world, the focus of this paper is to develop a novel and robust method to detect gender bias in corpora used to train LLMs in Spanish, a representative gendered language spoken by over 500 million people across diverse cultures and regions in the world. In Spanish, every noun carries a gender marker but only the words that refer to human beings are reflective of a potential gender bias.
@@
We propose a novel methodology that leverages state-of-the-art LLMs to detect and analyze gender bias in Spanish text corpora. Unlike traditional methods that rely on part-of-speech tagging or specific machine learning models to detect person references, our approach benefits from the contextual understanding capabilities of LLMs. By prompting LLMs to identify and classify nouns and pronouns by their reference to persons and their grammatical gender, we introduce a method that respects the linguistic structure of Spanish while providing the necessary depth of analysis.
@@
We first provide an overview of the most relevant literature on biases in NLP systems in English and non-English languages, followed by a detailed description of the proposed methodology. We empirically evaluate the proposed method on several benchmark datasets and uncover significant levels of gender imbalance with a male:female ratio ranging between 4:1 and 6:1, depending on the dataset. Interestingly, the English translations of the same corpora exhibit a male:female gender ratio of 1:1 to 3.5:1. Our results illustrate the importance of developing novel methods to detect gender bias in text corpora in gendered languages, such as Spanish. We conclude with a discussion, limitations, ethical implications, and lines of future work.
@@ 2 Related Work
There is a growing body of literature on biases in NLP systems, which has been summarized in a number of surveys (Blodgett et al., 2020; Stańczak and Augenstein, 2021; Bansal, 2022). In particular, Blodgett et al. (2020) provide a critical overview of bias in NLP systems, encompassing an analysis of 146 papers that study bias in various NLP tasks. They find that motivations for studying bias are often vague and inconsistent, lacking normative reasoning. For example, studies on embedding spaces (Bolukbasi et al., 2016; Caliskan et al., 2017; Gonen and Goldberg, 2019) illustrate biases but often do not critically engage with what constitutes bias. Similarly, works on language modeling (Lu et al., 2020; Bordia and Bowman, 2019), coreference resolution (Rudinger et al., 2018; Zhao et al., 2018), machine translation (Vanmassenhove et al., 2018; Stanovsky et al., 2019), and sentiment analysis (Kiritchenko and Mohammad, 2018) highlight bias in NLP but vary widely in their conceptualizations and methodologies. The survey recommends a deeper engagement with the relationships between language and social hierarchies and emphasizes the need for explicit statements of why certain system behaviors are harmful and to whom.
Recent advances in addressing biases in LLMs include various techniques focusing on detecting biases in the embeddings, prompt-based interactions, and contextual adaptations (Gallegos et al., 2024). However, we are not aware of any works aimed at the detection of bias in the training corpora, which is our focus.
Bias quantification approaches proposed to date have addressed three main areas: metrics for bias evaluation, datasets for bias evaluation, and methods leveraging generated text (Gallegos et al., 2024). Embedding-based metrics, such as the Sentence Encoder Association Test (May et al., 2019) and Contextualized Embedding Association Test (Guo and Caliskan, 2021), extend traditional bias metrics to contextualized embeddings. Probability-based metrics like pseudo-log-likelihood scores (Salazar et al., 2020) evaluate the likelihood of stereotypical versus anti-stereotypical sentences, while prompt-based metrics, including co-occurrence bias scores and sentiment analysis, measure bias in the generated text by comparing the distribution of terms associated with different social groups (Bordia and Bowman, 2019; Dhamala et al., 2021). These techniques provide comprehensive frameworks for assessing bias, ensuring a more nuanced and accurate understanding of how biases manifest in LLM outputs.
@@
From a language perspective, most existing research on bias in NLP has focused on English. For instance, Dhamala et al. (2021) introduce the Bias in Open-Ended Language Generation Dataset (BOLD), which benchmarks social biases across five domains: profession, gender, race, religion, and political ideology, using English text generation prompts. Similarly, Webster et al. (2018) address gender bias in coreference resolution with the Gendered Ambiguous Pronouns (GAP) corpus, a gender-balanced dataset of ambiguous pronoun-name pairs.
@@
While there is a body of literature on the impact of biases on various machine learning algorithms or studies related to stereotypes represented in datasets, to the best of our knowledge, there are no studies that have analyzed how to measure gender bias in texts written in gendered languages from the point of view of the prevalence of words referring to each gender. Addressing this gap is the focus of our research.
Our core contribution can be contextualized in a pyramid of methodologies to automatically detect gender bias in text corpora. Traditional word-counting methods would be located at the base of the pyramid. Unfortunately, their simplicity prevents their applicability to gendered languages like Spanish. Above this, we would find conventional NLP techniques, such as part-of-speech tagging. However, these methods might not reliably distinguish between gendered references of people versus animals or inanimate objects. At the apex of our methodological pyramid, we position our proposed LLM-based approach that not only identifies gendered words but also classifies them according to their role in conveying gender-related information within the text. Ultimately, we envision the use of these types of AI tools to automate bias evaluation with a comprehensive understanding of the underlying linguistic and social contexts.
This paper aims to shed light on the inherent challenges posed by gendered languages and propose a robust framework for detecting and measuring gender bias in Spanish text corpora. The insights derived from this study are expected to guide the development of fairer and more equitable language technologies, particularly in gendered languages.
@@ 3 Methodology
This section presents our methodology for assessing gender bias in language corpora. First, we describe a gender polarity method from the literature that has been proposed to measure gender-specific terms in English texts. We build upon this approach to propose a novel gender bias quantification method leveraging the LLMs’ natural language comprehension power to accommodate the complexities of gendered languages, using Spanish as a case study.
@@ 3.1 Gender Polarity
In a given text, the number of male tokens (denoted as $G_{M}$) and the number of female tokens ($G_{F}$) are counted. The gender bias in the text can then be expressed as the ratio $G_{M}:G_{F}$. The gender polarity method was designed for English texts, where gender differentiation in language usage is mostly captured through distinct pronouns and other gender-specific tokens. Nonetheless, this approach cannot be directly applied to gendered languages, such as Spanish, by merely translating the token lists to the target language. In gendered languages, specific parts of speech, such as nouns or adjectives, have inherent gender properties that often require different forms based on the gender of the subjects or objects they describe. This leads to a more complex interaction between grammatical gender and semantic content, necessitating a tailored method capable of accounting for the intrinsic gender agreement rules in such languages.
3.2 Measuring Gender Bias in Gendered Languages
We close the gap in bias quantification for gendered languages by proposing a method that takes inspiration from the gender polarity analysis and accommodates the specific grammatical and semantic features in gendered languages. We empirically evaluate the method on Spanish as an example of a gendered language.
@@ 3.2 Measuring Gender Bias in Gendered Languages
Note that we have consciously decided not to include adjectives, as they adapt their form to agree with the noun they modify, leading to redundancy in our analysis if those nouns are already accounted for. Furthermore, adjectives alone do not necessarily indicate the gender of the person being described unless paired directly with specific gendered nouns or pronouns, rendering the classification in step 2 impossible in many cases. Therefore, including adjectives could potentially negatively impact our method’s accuracy by introducing elements more reflective of grammatical structure rather than gender representation.
@@
3.3 Gender Bias Quantification Using LLMs
Implementing the pipeline described in the previous section using classical NLP methods would typically involve a combination of tools, leveraging part-of-speech tagging for step 1 and gender classification based, for instance, on word endings and dictionary lookups for step 3. However, accurately determining whether a noun or pronoun refers to a person (step 2) would require more sophisticated semantic analysis tools.
Given these challenges, we propose to leverage LLMs for their proficiency in understanding natural language nuances and context. LLMs can seamlessly integrate all steps of the proposed pipeline. After a process of manual interactive prompt-tuning, we designed a prompt that instructs the LLM to perform noun and pronoun identification, determine if these refer to human beings, and classify their grammatical gender, all in a single query. This approach utilizes the LLM’s ability to parse and interpret complex language structures and perform multiple tasks simultaneously.
@@ 3.3 Gender Bias Quantification Using LLMs
To analyze the gender representation in a given dataset, it is parsed sentence by sentence and each sentence is evaluated using the aforementioned prompt. We introduce the notation $L_{p,g}$, which denotes the number of words identified by our LLM-based approach, where $p\in\{P,N\}$ and $g\in\{M,F\}$ are binary variables that capture whether the word refers to a person ($P$) or not ($N$) and the word’s grammatical gender ($M$ for masculine, $F$ for feminine), respectively. Analogously to the gender polarity method described in Section 3.1, the gender bias in the given text is expressed by the ratio $L_{P,M}:L_{P,F}$.

### v2

# Leveraging Large Language Models to Measure Gender Representation Bias in Gendered Language Corpora

@@ Abstract
Gender bias in text corpora that are used for a variety of natural language processing (NLP) tasks, such as for training large language models (LLMs), can lead to the perpetuation and amplification of societal inequalities. This phenomenon is particularly pronounced in gendered languages like Spanish or French, where grammatical structures inherently encode gender, making the bias analysis more challenging. A first step in quantifying gender bias in text entails computing biases in gender representation, i.e., differences in the prevalence of words referring to males vs. females. Existing methods to measure gender representation bias in text corpora have mainly been proposed for English and do not generalize to gendered languages due to the intrinsic linguistic differences between English and gendered languages. This paper introduces a novel methodology that leverages the contextual understanding capabilities of LLMs to quantitatively measure gender representation bias in Spanish corpora. By utilizing LLMs to identify and classify gendered nouns and pronouns in relation to their reference to human entities, our approach provides a robust analysis of gender representation bias in gendered languages. We empirically validate our method on four widely-used benchmark datasets, uncovering significant gender prevalence disparities with a male-to-female ratio ranging from 4:1 to 6:1. These findings demonstrate the value of our methodology for bias quantification in gendered language corpora and suggest its application in NLP, contributing to the development of more equitable language technologies.
Leveraging Large Language Models to Measure Gender Representation Bias in Gendered Language Corpora
@@ 1 Introduction
Figure 1: Overview of the proposed method for the detection and measurement of biases in the prevalence of genders in gendered language corpora using LLMs.
@@
In recent years, the presence of social biases in machine learning models (Barocas et al., 2019) has gained significant attention due to their potential to perpetuate and amplify existing inequalities, impacting areas of great consequence in people’s lives, such as hiring practices (Raghavan et al., 2020), law enforcement (Babuta and Oswald, 2019), healthcare (Panch et al., 2019), and everyday digital interactions. Among various forms of bias, gender bias is particularly concerning because it affects roughly half of the global population and has pervasive effects across different sectors of society. This concern is amplified in the area of natural language processing (NLP), where biases present in training corpora can propagate through machine learning models and impact a broad range of applications (Kiritchenko and Mohammad, 2018; Costa-jussà and de Jorge, 2020).
@@
This paper focuses on representation bias – a type of bias where certain genders are underrepresented, stereotyped or portrayed in a limited manner – in the language corpora that are used to train large language models (LLMs). A core motivation for our study stems from the lifecycle of LLMs. From training to deployment, biases in the training data can not only be perpetuated but even amplified by the models (Kotek et al., 2023). The training corpora typically is obtained from diverse sources, including books, websites, social media, and other digital texts. If these sources contain biases – which they frequently do, reflecting societal prejudices and stereotypes – the models learn these biases as part of their training process. Hence, the first stage of data curation is of paramount importance. It involves selecting, cleaning, and annotating the training data to minimize biases as much as possible. Effective data curation requires an understanding of the types of biases that can exist and the methods to mitigate them, including removing or balancing biased content, ensuring diverse and representative data samples, and incorporating fairness and ethical guidelines in the data selection process (Gallegos et al., 2024).
In languages with grammatical gender (gendered languages), such as Spanish or French, this issue becomes even more complex. The grammatical structure may invisibly amplify certain gendered associations, impacting how often and in what context different genders are represented. In fact, research has shown that languages with grammatical gender exhibit more explicit gender biases due to frequent gender marking, which differs substantially from English, where gendered language use is mostly limited to pronouns and a few gender-specific terms (Kiritchenko and Mohammad, 2018; Costa-jussà and de Jorge, 2020). For example, masculine terms are often considered the ‘default’ in many gendered languages, which can implicitly prioritize male entities or perspectives. Studies indicate that grammatical gender affects model behavior in downstream tasks, including machine translation and sentiment analysis, leading to skewed model predictions that can disadvantage one gender over another (Gonen et al., 2019; Omrani Sabbaghi and Caliskan, 2022). For instance, Costa-jussà and de Jorge (Costa-jussà and de Jorge, 2020) demonstrate that machine translation systems tend to propagate gender biases more overtly in gendered languages. Consequently, adapting bias measurement methods to account for language-specific characteristics is essential for accurate assessments, as traditional methods used for English may not fully capture the nuances of gender bias in gendered languages like Spanish (Hellinger and Bußmann, 2001; Cho et al., 2021).
@@
Given the limitations of existing methods and the prevalence of gendered languages in the world, the focus of this paper is to develop a novel and robust method to detect gender representation bias in corpora used to train LLMs in Spanish, a representative gendered language spoken by over 500 million people across diverse cultures and regions in the world. In Spanish, every noun carries a gender marker but only the words that refer to human beings are reflective of a potential gender representation bias. To this end, we propose a novel methodology that leverages state-of-the-art LLMs to detect and analyze gender representation bias in Spanish text corpora. Unlike traditional methods that rely on part-of-speech tagging or specific machine learning models to detect person references, our approach benefits from the contextual understanding capabilities of LLMs. By prompting LLMs to identify and classify nouns and pronouns by their reference to persons and their grammatical gender, we introduce a method that respects the linguistic structure of Spanish while providing the necessary depth of analysis. Our methodology offers a proactive approach, enabling the identification of gender biases at the data level rather than addressing them post-training. By focusing on the unique grammatical structures of Spanish, we aim to provide a tool to help detect and mitigate biases early, contributing to the development of more equitable language models.
@@
We first provide an overview of the most relevant literature on biases in NLP systems in English and non-English languages, followed by a detailed description of the proposed methodology. We empirically evaluate the proposed method on several benchmark datasets and uncover significant levels of gender imbalance with a male:female ratio ranging between 4:1 and 6:1, depending on the dataset. Interestingly, the English translations of the same corpora exhibit a male:female gender ratio of 1:1 to 3.5:1. Our results illustrate the importance of developing novel methods to detect gender representation bias in text corpora in gendered languages, such as Spanish. We conclude with a discussion, limitations, ethical implications, and lines of future work.
@@ 2 Related Work
There is a growing body of literature on biases in NLP systems, which has been summarized in a number of surveys (Blodgett et al., 2020; Stańczak and Augenstein, 2021; Bansal, 2022). In particular, Blodgett et al. (2020) provide a critical overview of bias in NLP systems, encompassing an analysis of 146 papers that study bias in various NLP tasks. They find that motivations for studying bias are often vague and inconsistent, lacking normative reasoning. Works on language modeling (Lu et al., 2020; Bordia and Bowman, 2019), coreference resolution (Rudinger et al., 2018; Zhao et al., 2018), machine translation (Vanmassenhove et al., 2018; Stanovsky et al., 2019), and sentiment analysis (Kiritchenko and Mohammad, 2018) highlight bias in NLP but vary widely in their conceptualizations and methodologies. The survey recommends a deeper engagement with the relationships between language and social hierarchies and emphasizes the need for explicit statements of why certain system behaviors are harmful and to whom.
Gender imbalances in the training corpora are a significant and well-documented source of bias in machine learning. Numerous studies have shown that these imbalances can significantly influence model behavior. For example, Bolukbasi et al. (2016) found that word embeddings trained on biased data can amplify gender stereotypes in tasks like word similarity and analogical reasoning. Similarly, studies by Caliskan et al. (2017) and Brunet et al. (2019) demonstrate that biases present in training corpora can directly influence model outputs, perpetuating gendered stereotypes and imbalances in downstream tasks. Therefore, detecting and addressing gender imbalances in corpora is crucial to mitigating bias effects.
Recent advances in addressing biases in LLMs include various techniques focusing on detecting biases in the embeddings, prompt-based interactions, and contextual adaptations (Gallegos et al., 2024). However, we are not aware of any works aimed at the detection of gender representation bias in the training corpora, which is our focus.
Bias quantification approaches proposed to date have addressed three main areas: metrics for bias evaluation, datasets for bias evaluation, and methods leveraging generated text (Gallegos et al., 2024). Embedding-based metrics, such as the Sentence Encoder Association Test (May et al., 2019) and Contextualized Embedding Association Test (Guo and Caliskan, 2021), extend traditional bias metrics to contextualized embeddings. Probability-based metrics like pseudo-log-likelihood scores (Salazar et al., 2020) evaluate the likelihood of stereotypical versus anti-stereotypical sentences, while prompt-based metrics, including co-occurrence bias scores and sentiment analysis, measure bias in the generated text by comparing the distribution of terms associated with different social groups (Bordia and Bowman, 2019; Dhamala et al., 2021). These techniques provide comprehensive frameworks for assessing bias, enabling an accurate understanding of how biases manifest in LLM outputs.
@@
From a language perspective, most existing research about biases in NLP has focused on English. For instance, Dhamala et al. (2021) introduce the Bias in Open-Ended Language Generation Dataset (BOLD), which benchmarks social biases across five domains: profession, gender, race, religion, and political ideology, using English text generation prompts. Similarly, Webster et al. (2018) address gender bias in co-reference resolution with the Gendered Ambiguous Pronouns (GAP) corpus, a gender-balanced dataset of ambiguous pronoun-name pairs.
@@
While there is a body of literature on the impact of biases on various machine learning algorithms or studies related to stereotypes represented in datasets, to the best of our knowledge, there are no studies that have analyzed how to measure gender representation bias in texts written in gendered languages from the point of view of the prevalence of words referring to each gender. Addressing this gap is the focus of our research.
Our core contribution can be contextualized in a pyramid of methodologies to automatically detect gender representation bias in text corpora. Traditional male-female word-counting methods would be located at the base of the pyramid. Unfortunately, their simplicity prevents their applicability to gendered languages like Spanish. Above this, we would find conventional NLP techniques, such as part-of-speech tagging. However, these methods might not reliably distinguish between gendered references of people versus animals or inanimate objects. At the apex of our methodological pyramid, we position our proposed LLM-based approach that not only identifies gendered words but also classifies them according to their role in conveying gender-related information within the text. Ultimately, we envision the use of these types of AI tools to automate bias evaluation with a comprehensive understanding of the underlying linguistic and social contexts.
This paper aims to shed light on the inherent challenges posed by gendered languages and propose a robust framework for detecting and measuring gender bias in Spanish text corpora. The insights derived from this study are expected to guide the development of fairer and more equitable language technologies, particularly in gendered languages. While our primary focus is on detecting and quantifying gender representation bias, these capabilities represent a crucial first step in the debiasing process. By providing a reliable means to identify biases, our method enables the development of targeted interventions and ultimately supports creating more inclusive NLP applications.
@@ 3 Methodology
This section presents our methodology for assessing gender representation bias in language corpora. First, we describe a gender polarity method from the literature that has been proposed to measure gender-specific terms in English texts. We build upon this approach to propose a novel gender representation bias quantification method leveraging the LLMs’ natural language comprehension power to accommodate the complexities of gendered languages, using Spanish as a case study.
@@ 3.1 Gender Polarity
In a given text, the number of male tokens (denoted as $G_{M}$) and the number of female tokens ($G_{F}$) are counted. The gender representation bias in the text can then be expressed as the ratio $G_{M}:G_{F}$. The gender polarity method was designed for English texts, where gender differentiation in language usage is mostly captured through distinct pronouns and other gender-specific tokens. Nonetheless, this approach cannot be directly applied to gendered languages, such as Spanish, by merely translating the token lists to the target language. In gendered languages, specific parts of speech, such as nouns or adjectives, have inherent gender properties that often require different forms based on the gender of the subjects or objects they describe. This leads to a more complex interaction between grammatical gender and semantic content, necessitating a tailored method capable of accounting for the intrinsic gender agreement rules in such languages.
3.2 Measuring Gender Representation Bias in Gendered Languages
We close the gap in the quantification of gender representation bias in gendered languages by proposing a method that takes inspiration from the gender polarity analysis and accommodates the specific grammatical and semantic features in gendered languages. We empirically evaluate the method on Spanish as an example of a gendered language.
@@ 3.2 Measuring Gender Representation Bias in Gendered Languages
Note that we have consciously decided not to include articles, determinants and adjectives, as they adapt their form to agree with the noun they modify, leading to redundancy in our analysis if those nouns are already accounted for.
@@
An important consideration in analyzing Spanish is the traditional convention of using the male plural form to refer to groups that may include both men and women (e.g., doctores for doctors, including both male and female doctors). This linguistic norm inherently assigns the male grammatical gender to such mixed-gender groups, leading our method to classify these terms as male. This convention, although prevalent in many gendered languages, can contribute to the under-representation of females. Currently, in Spanish as in other gendered languages, listing explicitly both genders is the preferred form and has become the new standard [^1] https://www.unwomen.org/sites/default/files/Headquarters/Attachments/Sections/Library/Gender-inclusive%20language/Guidelines-on-gender-inclusive-language-es.pdf (e.g., doctoras y doctores collectively referring to female and male doctors). Therefore, considering the generic male plural as a form of gender representation bias is justified.
3.3 Gender Representation Bias Quantification Using LLMs
Implementing the pipeline described in the previous section by means of classical NLP methods would typically involve a combination of tools, leveraging part-of-speech tagging for step 1 and gender classification based, for instance, on word endings and dictionary lookups for step 3. However, accurately determining whether a noun or pronoun refers to a person (step 2) would require more sophisticated semantic analysis tools.
Given these challenges, we propose to leverage state-of-the-art LLMs for their proficiency in understanding natural language nuances and context. LLMs can seamlessly integrate all steps of the proposed pipeline. After a process of manual interactive prompt-tuning, we designed a prompt that instructs the LLM to perform noun and pronoun identification, determine if these refer to human beings, and classify their grammatical gender, all in a single query. This approach utilizes the LLM’s ability to parse and interpret complex language structures and perform multiple tasks simultaneously.
An important advantage of our method is its scalability to other gendered languages beyond Spanish. The use of state-of-the-art multilingual large language models enables our approach to handle a wide range of languages that incorporate grammatical gender. These models have been trained on diverse linguistic data, allowing them to recognize and analyze gendered terms in languages with varying structures and grammatical rules.
@@ 3.3 Gender Representation Bias Quantification Using LLMs
To analyze the gender representation in a given dataset, it is parsed sentence by sentence and each sentence is evaluated using the aforementioned prompt. We introduce the notation $L_{p,g}$, which denotes the number of words identified by our LLM-based approach, where $p\in\{P,N\}$ and $g\in\{M,F\}$ are binary variables that capture whether the word refers to a person ($P$) or not ($N$) and the word’s grammatical gender ($M$ for masculine, $F$ for feminine), respectively. Analogously to the gender polarity method described in Section 3.1, the gender representation bias in the given text is expressed by the ratio $L_{P,M}:L_{P,F}$.